OpenStudy (anonymous):

How do you find the point of intersection(s) for x = 2y2 + 3y + 1 and 2x + 3y2 = 0 A) You cannot find points of intersections for non-functions. B) Plug in 0 for x into both equations and solve for y. Then plug that answer back into the other equation to find the corresponding x-coordinate. C) Solve both equations for x and set them equal to each other. This will give you the y-coordinates of the points of intersection. Then plug back into one of the equations to find the corresponding x-coordinates. D) Solve both equations for y and set them equal to each other. This will giv

OpenStudy (dopeness_):

You use method C, or method D. They are both technically the same method. One you find x first, the other you find y first. (to say to the power of, we write ^(number)...) x = 2y^2 + 3y + 1 and 2x + 3y2 =0 x=-3y^2/2 2y^2+3y+1=-3y^2/2 4y^2+6y+2=-3y^2 7y^2+6y+2=0 You have a quadratic. Solve it. You'll get TWO y values (this means the curves intersect at two places). Find the two y values, then sub BOTH y values (at different times) into the original formula and you'll get the corresponding x values.